Suppose you want to introduce a third grader to the standard algorithm for computing 181 + 331. Explain how to find the sum with base-ten blocks following the same sequence of steps as the standard algorithm.
1
Expert's answer
2012-07-05T07:53:34-0400
181 + 331

First, getting the sum of lowest block:
1 + 1 = 2
2 < 10, so:
Second, getting the sum of the next block:
8 + 3 = 11
11 > 10, so: 11 - 10 = 1 stay in this block and 10 / 10 = 1 will pass to the next block.
Thie rd, getting the sum of the highest block with the passed 1:
1 + 3 + 1 = 5
5 < 10, so the result is 512.